I get so confused with these problems I just don't understand.
It says Solve the Quadratic equations by factoring
(2x-7)(7x+2)=0 And so how would I go about solving this?
turn it into 14x(squared)-45x-14=0??? and then? ugggghh
Some one please help me!!
Thank you for yout time.

Answer by checkley75(3666)

(2x-7)(7x+2)=0 there are two ways to go:
---------------------------------------------------------------
solve each of the factors for x thus:
2x-7=0, 2x=7, x=7/2, x=3.5 answer.
7x+2=0, 7x=2, x=2/7 answer.
---------------------------------------------------------------
OR using the FOIL (first, outer, inner, last) multiplication process we get:
14x^2+4x-49x-14=0
14x^2-45x-14=0